OpenStudy (anonymous):

Grady is comparing three investment accounts offering different rates. Account A: APR of 4.95% compounding monthly Account B: APR of 4.85% compounding quarterly Account C: APR of 4.75% compounding daily Which account will give Grady at least a 5% annual yield? Account A Account B Account C Account B and Account C

OpenStudy (tkhunny):

Well, do it! \(\left(1 + \dfrac{0.0495}{12}\right)^{12} - 1\) \(\left(1 + \dfrac{0.0485}{4}\right)^{4} - 1\) \(\left(1 + \dfrac{0.0475}{365}\right)^{365} - 1\)

OpenStudy (anonymous):

Account A!

OpenStudy (tkhunny):

Did you do them all or just decide that the first one that worked was enough?

OpenStudy (anonymous):

No, I did them all. I got 0.050 for A, 0.0493 for B, and 0.0486 for account C

OpenStudy (tkhunny):

Perfect!

OpenStudy (anonymous):

Thank you!